OpenStudy (anonymous):

Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Provide the solution to the system. 14x – y = –10 7x + 2y = 15

OpenStudy (anonymous):

Solve for x, and then subsitute the value of x into the equation.

OpenStudy (anonymous):

Add y to the right side then divide the right side by 14.

OpenStudy (anonymous):

You should get x=y-5/7

OpenStudy (anonymous):

Plug that into the second equation.

OpenStudy (wolf1728):

A) 14x – y = –10 B) 7x + 2y = 15 Multiply equation A by 2. Add this to equation B. You now have an equation solved for x and now you you can solve for y.

OpenStudy (anonymous):

Thanks!

OpenStudy (wolf1728):

u r welcome :-)